WARNING: This web is not used anymore. Please use PDBService.MinuteS12April07 instead!

Minutes 12 April 07

  • Phone:

    • FNAL: Eric
    • NDGF: Olli
    • Triumf: Jim Li

  • CERN:
    • 3D: Dirk, Eva
    • POOL: Ioannis
    • CORAL: Radovan
    • ATLAS: Hans
    • Taiwan: Dave

  • Sites status:
    • BNL:
      • new RAC database available outside the firewall
      • Streams throughput tests: 4000 LCRs per second
      • Working on the agent configuration – point Dirk Geppert about the timezone configuration problem
    • RAL:
      • Intervention today, completed successfully
    • PIC:
      • second cluster for ATLAS installation ongoing
    • SARA:
      • Dirk will check with Alexander when the new cluster will be available
    • Taiwan:
      • Performance tests ongoing
    • Triumf:
      • Eva has run new tests, some of them have achieved up to 5000 LCRs per second
      • No changes on Triumf side
      • Run iperf tests together with streams tests to determine network bandwidth
      • Bandwidth monitoring in order to detemine if the bandwidth is the limitation * NDGF:
      • Nothing to report
      • Timeline for the cluster installation to be provided
  • FTS:
    • Installation at PIC run in problems: archive log area full, no RMAN backups configured
    • Sites should run monitoring over FTS databases (similar as they do with the experiments databases)
    • Dirk will send email to collect information about the databases running by the sites (FTS, Castor, ...)

  • Sites must announce the interventions on the general email list , egee broadcast, ...

  • Experiments status:
    • ATLAS:
      • COOL conditions replication is the priority
      • TAGs will be replicated by file, not using Streams
      • Tests starting on 15th April
      • After several weeks of stable running, then the recovery exercise will be performed with the sites participation (aprox. middle of May)
      • Streams synchronization at Tier0 exercise to be performed as well
      • ATLAS Scalability tests: CNAF, LYON and Stefan have been contacted
    • LHCb:
      • Check with Marco about the starting date

  • Frontier CMS:
    • Tracking squid details for the monitoring
    • Dirk will check the milestones for CMS with Lee
Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2007-04-13 - EvaDafonte
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2022 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback